# Lessons from Learning to Spin “Pens”

### v1

@@ 1 Introduction
Pen spinning has been studied from several perspectives. Classic robotics works demonstrate rotating wooden blocks with open-loop force control [1]. With high-speed cameras and advanced hardware, agile pen spinning can also be achieved [6]. However, these methods rely on accurate object models and cannot generalize to unseen objects. On the other hand, learning-based methods hold the promise of being generalizable with large-scale data. They have indeed achieved significant progress either with imitation learning [7, 8, 9] or sim-to-real [3, 10, 11, 12]. However, they have only demonstrated manipulation of regular spherical or cuboid-shaped objects, and none can extend the capability to pen-like objects. We attribute this to two reasons: For imitation learning, current teleoperation systems fail at collecting complex and dynamic demonstrations such as pen spinning; for sim-to-real, bridging the gap for dynamic tasks becomes substantially difficult.
@@ 2 Related Work
Classic in-hand manipulation. In-hand manipulation has been studied for decades [2, 16]. Classical methods rely on an accurate model and analytically plan a sequence of motions to control the object. For example, Han and Trinkle [17] manipulate objects using sliding, rolling, and finger gaiting motions, while Bai and Liu [18] studies the collaboration of fingers and the palm. Open-loop manipulation also shows surprising robustness and dexterous behavior [14, 15].  Sieler and Brock [19] uses linearized feedback-control for in-hand manipulation with a soft hand. State-of-the-art systems in this category include full $SO(3)$ reorientation using a compliance-enabled hand [20] and an accurate pose tracker [21]. However, most methods cannot manipulate pen-like objects due to their complex and dynamic nature. Extrinsic dexterity [22] can also be used to achieve dynamic manipulation, but a precise model is necessary. In contrast, our method uses human priors to build a simulator environment but does not rely on an accurate model during deployment.
@@ 4.3 Sensorimotor Policy Training
Q6: How do pre-training and fine-tuning contribute to the final performance? Our approach is first pre-trained in simulation and then fine-tuned using real-world data. We study the contribution of each part in Table 2. With only pre-training, the policy has limited effectiveness in the real world. It rarely completes finger gaiting on Objects C and D, and the success rate is also low for the remaining objects. This is mainly because the physics gap between simulation and reality becomes more significant in our task. With only behavior cloning, the approach also does not perform well. It has a 50% lower rotation radius on Training Objects. On out-of-distribution objects, the success rate drops to less than 20%, indicating that its generalization capability is limited. This is because we only have 15 trajectories for each object, and the policy tends to overfit to that data.
@@ Acknowledgments
Xiaolong Wang’s lab is supported, in part, by Amazon Research Award, Intel Rising Star Faculty Award, and Qualcomm Innovation Fellowship. Haozhi Qi and Jitendra Malik are supported in part by ONR MURI N0001421-1-2801. Haozhi Qi and Yi Ma are supported in part by ONR N00014-22-1-2102.

### v2

@@ 1 Introduction
Pen spinning has been studied from several perspectives. Classic robotics works demonstrate rotating wooden blocks with open-loop force control [1]. With high-speed cameras and advanced hardware, agile pen spinning can also be achieved [6]. However, these methods rely on accurate object models and cannot generalize to unseen objects. On the other hand, learning-based methods hold the promise of being generalizable with large-scale data. They have indeed achieved significant progress either with imitation learning [7, 8, 9] or sim-to-real [3, 10, 11, 12]. However, they have only demonstrated manipulation of regular spherical or cuboid-shaped objects, and none can extend the capability to pen-like objects. We attribute this to two reasons: For teleoperation-imitation pipeline, current teleoperation systems fail at collecting complex and dynamic demonstrations; for sim-to-real, bridging the gap for dynamic tasks becomes substantially difficult.
@@ 2 Related Work
Classic in-hand manipulation. In-hand manipulation has been studied for decades [2, 16]. Classical methods rely on an accurate model and analytically plan a sequence of motions to control the object. For example, Han and Trinkle [17] manipulate objects using sliding, rolling, and finger gaiting motions, while Bai and Liu [18] studies the collaboration of fingers and the palm. Mordatch et al. [19] demonstrates object rotation in simulation by trajectories optimization. Li et al. [20] learns a object-level impedance controller for both grasping and rotating objects. Open-loop manipulation also shows surprising robustness and dexterous behavior [14, 15].  Sieler and Brock [21] uses linearized feedback-control for in-hand manipulation with a soft hand. State-of-the-art systems in this category include full $SO(3)$ reorientation using a compliance-enabled hand [22] and an accurate pose tracker [23]. However, most methods cannot manipulate pen-like objects due to their complex and dynamic nature. Extrinsic dexterity [24] can also be used to achieve dynamic manipulation, but a precise model is necessary. In contrast, our method uses human priors to build a simulator environment but does not rely on an accurate model during deployment.
@@ 4.3 Sensorimotor Policy Training
In the above experiments the initial hand and object configurations are chosen from the replay trajectory dataset. This setting brings advantage to the oracle replay baseline. To comprehensively study the performance, we also conduct real-world experiments where the objects are initialized at a randomly chosen stable grasp. We choose 10 random grasps, do 5 trials for each. For each random grasp, we generate the oracle replay trajectory by running the oracle policy in the simulator and select the best trajectory among 1000 simulated environments. In this setting, our method achieves 22.0% (from 54% to 78.0%) better success rate on object D, 36% (from 46 to 82) on object E, and 40% (from 34% to 74%) on object F. The reason that our policy does not drop compared to the number if Table 1 is because it has seen much more diverse data via simulation pre-training.
@@
Q6: How do pre-training and fine-tuning contribute to the final performance? Our approach is first pre-trained in simulation and then fine-tuned using real-world data. We study the contribution of each part in Table 2. With only pre-training, the policy has limited effectiveness in the real world. It rarely completes finger gaiting on Objects C and D, and the success rate is also low for the remaining objects. This is mainly because the physics gap between simulation and reality becomes more significant in our task. With only behavior cloning, the approach also does not perform well. For this experiment, we use ACT [63] as the architecture, which is one of the best imitation learning algorithms. It has a 50% lower rotation radius on Training Objects. On out-of-distribution objects, the success rate drops to less than 20%, indicating that its generalization capability is limited. This is because we only have 15 trajectories for each object, and the policy tends to overfit to that data.
@@ 5 Conclusion and Lessons
In addition, our work assumes the object is placed at a stable grasp position following previous work [11, 10]. Incorporating more advanced grasping work [64] or consider chaining different skills [45] together would make our work more general.
@@ Acknowledgments
Xiaolong Wang’s lab is supported, in part, by Amazon Research Award, Intel Rising Star Faculty Award, Qualcomm Innovation Fellowship, and gifts from Meta. Haozhi Qi and Jitendra Malik are supported in part by ONR MURI N0001421-1-2801. Haozhi Qi and Yi Ma are supported in part by ONR N00014-22-1-2102.
@@ A.1 Training Hyper-parameters
A.2 Domain Randomization Parameters
The domain randomization list is as follows:
Parameter Range
Object Scale x[0.95, 1.05]
[…]
Coefficient of Friction (obj and fingertip) [0.3, 3.0]
External Disturbance (0.2, 0.25)
PD Controller Stiffness [2.5, 3.5]
PD Controller Damping [0.09, 0.11]
Observation Noise N(0, 0.02) (rad)
Action Noise N(0, 0.01) (rad)
Table 8: Domain Randomization Parameters. The object scale range is multiplied by the original scale. Observation and action randomizations follow a Gaussian distribution with the specified radius. Other randomizations are uniformly sampled from the specified range. Following [3], we apply a random disturbance force to the object during training whose scale is $0.2m$ with probability 0.25 where $m$ is the object mass.
A.3 System Identification
We tune the P and D gain in the low level controller according to the following two metrics: 1) Before we do data collection, we first replaying several finger gait trajectory without objects in-hand. We do this for both sim and real and compare the errors between joint positions. We try to minimize the error by tuning the P and D gains simultaneously in sim and real. 2) We also command the sin and cos waves of each joint and observe the errors between sim and real. We also include action noises during training so that the policy can be robust to real-world actuator noises.
A.4 Physical Properties of Real-world Objects
The physical properties of real-world objects are shown in Table 9.
Obj ID Mass (g) Length (cm) Contact Part Diameter (mm)
A 16.7 22.21 34.66
B 10.8 14.41 27.34
C 21.4 17.43 34.73
D 29.6 18.30 35.50
E 32.4 19.12 29.73
F 36.3 19.00 31.17
G 22.3 20.09 36.88
H 21.5 14.89 31.81
I 26.0 15.00 37.93
J 49.7 15.75 35.80
Table 9: Physical Parameters of Real-world Objects. We show the object’s mass, length, and the diameter of contact parts. The mass distribution ranges from 10g to 50g, length varies from 15 cm to 22 cm, and diameter ranges from 29 mm to 34 mm. The materials used include 3D-printed components, leatheroid, and various plastics, all of which have different frictional characteristics.